Solve the IVP using the method of Laplace transforms AND one other method of your choice. y +5y +6y= 2e ; y(0)=1, y(0) = 3

Detailed answer with another method then the Laplace transforms
Answer #1

2e-t or Given IVP is 9+ 5y +6y=zet. First we solve * by using variation of parameters. Comparing with 9+ pylt Qy=R we haveIt) = f und e+ 2e-t t - е. et —3t - at t + = -e seztat --eat s det = sest set alt =pſ et dt = zet Yplt) = u. flt) + Vig1+) -& - se+95+10 = A(5+2)(5+3) + B (5+1)(5+3) 76(5+1)(5+2) -# fubstitute s=-1 in # we get A.1.2 =) A=1 substitute s=-2 in # we ge